OpenStudy (anonymous):

Write the equation of a circle with a center at (-3, -4) and a radius of 3.

OpenStudy (anonymous):

\[(x+3)^{2}+ (y+4)^{2}= 9\]

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2, plug in the given. (h,k) points r radius.. x^2+y^2+6x+8y+16=0 I think..